OpenStudy (anonymous):

Use identities (no calculators) to find the exact value for (sin 38)(sin 112) -( cos 38)(cos 112).

OpenStudy (anonymous):

same trick as before. this is the addition angle formula for cosine{ \[cos(a+b)=cos(a)cos(b)-sin(a)sin(b)\]

OpenStudy (anonymous):

here \[a=38,b=112, a+b=150\]

OpenStudy (anonymous):

well on careful inspection i see that this is it's negative, but that is ok \[cos(150)=-\frac{\sqrt{3}}{2}\] so your answer here is \[\frac{\sqrt{3}}{2}\]

OpenStudy (anonymous):

Why does the negative turn to a positive?

OpenStudy (anonymous):

oh because the "addition angle" formula is not the one written in the question. it is its negative